show me how to solve -85.68 + delta = 0 Optien Drieing Mdel cuse Ageroadh Pert talio AAdbitase 61T U-e-e 100 1-22/403 T-I y 4- dundens Coll Alarak Cpton Slack Ina C1

how to solve for delta step by step pls hpw did they get 65.3491
Answer #1

Steps to find delta

1) Calculate upward and downward movement represented by u and d from the given values

= 1221403 d = 1 = 0.818734

2) Find the stock price if the stock worth 100 moves corresponding to d and u values.

Stock 7100x1.22 1703 y Soxd = coox o818731 = 818731

3) Find the option values in case of upward or downward movement. Here the exercise price is 90. Hence, the value of option is Maximum of (Spot price- Exercise price, 0) which means it is a call option. ( A call option will be exercised only if the spot price of the share underlying is higher than the exercise price)

Call Ootion top max (So U-ko) 2 max (122.1403 - 90, o) 32. 03 tdown-max Sod-ko) max (51.8731-90,0) witht man

4) As per Arbitrage Portfolio Theory the following functions must return the same value

Arbitrage Portfolio totup ta Sao YU 732.1403+A. 122.1403 1 J soloe for delta A -tdow + A sod 6.349) - 0 818731

5) By solving the above equations, we can find Delta ie: -fup + delta x Sou = -fdw + delta x Sod which when solved will give the equation for delta.

1 = I tup - Sol - f down Sod 32./103 - 0 ( 122 1403-81.87 31 LA = 0.798176

6) Now, substituting the delta value in the functions as per (4) will return you the value of 65.3491 :

-fup + delta x Sou= -32.1403+ 0.798176 x 122.1403= 65.3491

-fdw + delta x Sod = -0+ 0.798176 x 81.8731= 65.3491

7) Find the option price now

62.7867 to A So-IPU to = 6.798176x100-62.7867 to = 17.0309 = 65,3491er = 65.8491e = 62,7867 -0.04